# GARCH copulas and GARCH-mimicking copulas

### v1

@@ 3.1 ARCH and GARCH processes
    A process of ARCH(1) type is a process of GARCH(1,1) type where the volatility function $\sigma(x,s)$ depends only on $x$.
@@
    The classical GARCH(1,1) process is a process of GARCH(1,1) type type in which the volatility function satisfies
@@
    A process of GARCH(1,1) type is said to have symmetric distribution, if the distribution of $\epsilon_{1}$ has a density that is symmetric around zero, and to have symmetric volatility if $\sigma(x,w)$ is an even function of $x$. It is said to be symmetric if it has symmetric distribution and symmetric volatility.
@@
The mathematical properties of the classical GARCH(1,1) process have been widely studied and the conditions on the parameters in part 3 of Definition 3 are known to guarantee a strictly stationary solution to the equations (4(Nelson,, 1990; Bougerol and Picard,, 1992). For a finite-variance solution it is usual to assume the stronger condition that $\alpha_{1}+\beta_{1}<1$. It is known that the marginal or stationary distribution $F_{X}$ has regularly varying (power-law) tails and Mikosch and Stărică, (2000) showed that the tail index $\zeta$ may be calculated as the root of the equation $\mathbb{E}\left(\left(\alpha_{1}\epsilon_{1}^{2}+\beta_{1}\right)^{\zeta/2}\right)=1$. However, the exact functional forms of $F_{X}$ or its density $f_{X}$ are not available.
@@ 3.2 The first-order copula of GARCH(1,1)-type processes
Simple calculations show that the joint density of a pair of successive variables $(X_{1},X_{2})^{\top}$ in a process of GARCH(1,1) type is given by
@@
where $f_{\sigma}$ is the marginal density of the volatility process $(\sigma_{t})_{t\in\mathbb{Z}}$ and $f_{\epsilon}$ is the density of the innovations. Applying (3), the density $c_{1}$ of the copula $C_{1}$ of $(X_{1},X_{2})^{\top}$ is thus
@@
The key insight from these calculations is the following.
@@
    An even volatility function also implies that $f_{X_{1},X_{2}}(x,y)=f_{X_{1},X_{2}}(-x,y)$ and so $(X_{1},X_{2})^{\top}$ is v-symmetric about zero and hence radially and jointly symmetric about the origin.
In case (i) above we can apply Lemma 1 to infer that the copula $C_{1}$ is h-symmetric and we can apply Proposition 3 to infer that its density may be written in the form $c_{1}(u,v)=c_{X_{1},|X_{2}|}(u,|2v-1|)$ where $c_{X_{1},|X_{2}|}$ is the density of the copula of $(X_{1},|X_{2}|)^{\top}$. In case (ii) we infer that the copula $C_{1}$ is jointly symmetric and its density may be written in the form $c_{1}(u,v)=c_{|X_{1}|,|X_{2}|}(|2u-1|,|2v-1|)$ where $c_{|X_{1}|,|X_{2}|}$ is the density of the copula of $(|X_{1}|,|X_{2}|)^{\top}$.
Despite the strong symmetry revealed by Proposition 4, there is one important form of symmetry that symmetric processes of GARCH(1,1) type do not have in general, namely exchangeability. It is not true that $f_{X_{1},X_{2}}(x,y)=f_{X_{1},X_{2}}(y,x)$ for all $(x,y)^{\top}\in\mathbb{R}^{2}$ so the bivariate distributions of the pairs $(X_{1},X_{2})^{\top}$ and $(X_{2},X_{1})^{\top}$ differ. This implies that ARCH and GARCH are not time-reversible models. In general $f_{X_{2}\mid X_{1}}(y\mid x)\neq f_{X_{1}\mid X_{2}}(y\mid x)$ and thus, in an intuitive sense, they are models in which the predictive distribution of the future given the past differs from the “predictive distribution” of the past given the future. From an econometric point of view, this would be considered natural. The non-exchangeability carries over to the copula so that we have in general that $C_{1}(u,v)\neq C_{1}(v,u)$ on $[0,1]^{2}$. Despite this observation, it is notable that the majority of the copulas employed in vine copula models of financial time series are exchangeable ones.
@@ 3.3 Computational examples
In this section we use computational methods to approximate and plot the joint densities $f_{X_{1},X_{2}}$ and copula densities $c_{1}$ in a number of models of GARCH(1,1) type. We concentrate on processes of ARCH(1) type in our illustrations since their first-order Markov structure simplifies computation of the unknown marginal density $f_{X}$ and distribution function $F_{X}$, which are key ingredients in (9). Beginning with the equation
@@
We also consider the classical GARCH(1,1) model with Gaussian innovations. To analyse this process we use simulation and estimation to find a parametric model for the unknown marginal density $f_{\sigma}$ of the volatility process. We can then integrate out volatility in expressions like (6), (7) and (8) [^1] All details of the model for $f_{\sigma}$ and the discretization method for ARCH(1)-type processes can be found in the code that accompanies this paper..
The parameter values for all the models we consider are summarized in Table 1.
@@
Table 1: Parameter values for processes analysed in paper. The volatility function $\sigma(x,s)$ in all models satisfies $\sigma(x,s)^{2}=\alpha_{1}+(\alpha_{1}+\gamma_{1}I_{\{x<0\}})x^{2}+\beta_{1}s^{2}$. Innovation distributions used are either Gaussian, Student t with $\nu=4$ or $\nu=2.5$ degrees of freedom or skewed Student t with $\nu=4$ degrees of freedom and skewness parameter $\lambda=0.8$; all innovation distributions are scaled to have mean zero and variance one.
Figure 1: Contour plots of $f_{X_{1},X_{2}}(x_{1},x_{2})$ for symmetric GARCH(1,1)-type processes; see Table 1 for all parameter values. Top left, top right, bottom left: classical ARCH(1) processes with, respectively, Gaussian, Student t4 and Student t2.5 innovations. Bottom right: classical GARCH(1,1) process with Gaussian innovations.
@@
  1. 1.
    Classical ARCH(1) process with Gaussian innovations (symmetric) and parameter values $\alpha_{0}=0.4$ and $\alpha_{1}=0.6$ as in Table 1; this process has unit variance and tail index $\zeta=3.82$. The joint density $f_{X_{1},X_{2}}$ and copula density $c_{1}$ are shown in the top left panels of Figures 1 and 2 respectively. The non-exchangeability, evidenced by a lack of symmetry in the dashed lines $x_{2}=x_{1}$ and $v=u$, is evident in both pictures. We can also observe the joint symmetry, that is the rotational symmetry of order 2 as well as the h-symmetry and v-symmetry.
    Changing $\alpha_{1}$ changes the degree of dependence in the pictures. In particular, as $\alpha_{1}\to 0$ the copula converges to the independence copula while increasing $\alpha_{1}$ serves to strengthen the cross-shaped character of the copula density. For normal innovations, the theoretical maximum value of $\alpha_{1}$ to obtain a strictly stationary process is approximately 3.56, but for values $\alpha_{1}\geqslant 1$ we obtain infinite-variance processes which are seldom used in practical applications and become increasingly pathological with increasing $\alpha_{1}$ as their tail indexes converge to zero (Mikosch and Stărică,, 2000).
  2. 2.
@@
    Classical ARCH(1) process with Student innovations (symmetric). This is shown in the top right panels of Figures 1 and 2 for a process with the same parameters as in the previous example and degree of freedom $\nu=4$ for the innovation distribution; the tail index of the process is 2.76. Remarkably, the resulting joint density appears to be very close to spherical and the copula appears to be very close to an exchangeable copula for this combination of parameters; the joint symmetry is again apparent. The bottom left panels of Figures 1 and 2 show a process in which $\nu=2.5$ and the tail index is 2.09; these show non-exchangeability more clearly.
@@
    It is notable that, although the tail index of the marginal distribution increases with decreasing degree of freedom (increasing tail index of the innovation density), the dependence actually weakens. We can estimate a distance $D_{1}=\int_{0}^{1}\int_{0}^{1}|c_{1}(u,v)-1|\mathrm{d}u\mathrm{d}v$ between the copula density $c_{1}$ and the density of the independence copula; for the Gaussian, t4 and t2.5 copulas we obtain respectively 0.213, 0.158 and 0.092.
  3. 3.
    Classical GARCH(1,1) process with Gaussian innovations (symmetric). This is shown in the bottom right panels of Figures 1 and 2 for a process with Gaussian innovations and parameters as in Table 1, which has variance one and tail index 4.09.

### v2

@@ 2.1 ARCH and GARCH processes
    A process of ARCH(1) type is a process of GARCH(1,1) type where $\sigma(x,s)$ depends only on $x$.
@@
    The classical GARCH(1,1) process is a process of GARCH(1,1) type in which $\sigma(x,s)$ satisfies
@@
    A process of GARCH(1,1) type is said to have symmetric distribution, if the distribution of $\epsilon_{1}$ has a density that is symmetric around zero, and to have symmetric volatility if $\sigma(x,s)$ is an even function of $x$. It is said to be symmetric if it has symmetric distribution and symmetric volatility.
@@
The mathematical properties of the classical GARCH(1,1) process have been widely studied and the conditions on the parameters in part 3 of Definition 1 are known to guarantee a strictly stationary solution to the equations (2(Nelson,, 1990; Bougerol and Picard,, 1992). For a finite-variance solution it is usual to assume the stronger condition that $\alpha_{1}+\beta_{1}<1$. It is known that the marginal or stationary distribution $F_{X}$ has regularly varying (power-law) tails with a tail index $\zeta$ that may be calculated as the root of the equation $\mathbb{E}\left(\left(\alpha_{1}\epsilon_{1}^{2}+\beta_{1}\right)^{\zeta/2}\right)=1$ (Mikosch and Stărică,, 2000). However, the exact functional forms of $F_{X}$ and its density $f_{X}$ are not available.
2.2 Symmetry concepts for bivariate distributions
In our analysis of the pair copulas of GARCH processes, we first consider models with symmetric innovation distributions. To describe the resulting copulas, it helps to recall the definitions of radial and joint symmetry in Nelsen, (2006) and to identify two further concepts we call h- and v-symmetry (symmetry about horizontal and vertical axes).
Definition 2.
Let $(Y,Z)^{\top}$ be a random vector taking values in a set $A\subseteq\mathbb{R}^{2}$.
  1. (i)
    If $(Y,Z-b)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(Y,b-Z)^{\top}$ for some $b\in\mathbb{R}$ we say that $(Y,Z)^{\top}$ is h-symmetric about $b$.
  2. (ii)
    If $(Y-a,Z)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(a-Y,Z)^{\top}$ for some $a\in\mathbb{R}$ we say that $(Y,Z)^{\top}$ is v-symmetric about $a$.
  3. (iii)
    If $(Y-a,Z-b)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(a-Y,b-Z)^{\top}$ for $a,b\in\mathbb{R}$ we say that $(Y,Z)^{\top}$ is radially symmetric about $(a,b)$.
  4. (iv)
    If (i), (ii) and (iii) all hold, we say that $(Y,Z)^{\top}$ is jointly symmetric about $(a,b)$.
Remarks 1.
  1. 1.
    We also describe the joint distribution function (df) $F_{YZ}$ of $(Y,Z)^{\top}$ as having these symmetries.
  2. 2.
    If (i) holds, the marginal distribution of $Z$ is symmetric about $b$; if (ii) holds, the marginal distribution of $Y$ is symmetric about $a$. The reverse implications are not true.
  3. 3.
    H-symmetry and v-symmetry together imply radial and joint symmetry. However radial symmetry does not necessarily imply h-symmetry and v-symmetry, although it does imply symmetry of both marginal distributions.
  4. 4.
    Nelsen, (2006) notes that jointly symmetric distributions have zero correlation (when finite second moments permit the measure to be defined). In Section S1 of the Online Supplement, we show that h-symmetry (or v-symmetry) on its own is sufficient for an absence of both correlation and rank correlation. Despite the absence of correlation, distributions with these forms of symmetry can have quite strong forms of dependence.
If the distribution functions $F_{Y}$ and $F_{Z}$ of $Y$ and $Z$ are continuous, the random vector $(U,V)^{\top}=(F_{Y}(Y),F_{Z}(Z))^{\top}$ has distribution function $C$, known as the copula of $(Y,Z)^{\top}$, and the symmetries have the following implications for $C$.
Lemma 1.
  1. (i)
    If $(Y,Z)^{\top}$ is h-symmetric then C(u,v)=uC(u,1v)C(u,v)=u-C(u,1-v), for $(u,v)\in[0,1]^{2}$.
  2. (ii)
    If $(Y,Z)^{\top}$ is v-symmetric then C(u,v)=vC(1u,v)C(u,v)=v-C(1-u,v), for $(u,v)\in[0,1]^{2}$,
  3. (iii)
    If $(Y,Z)^{\top}$ is radially symmetric then C(u,v)=u+v+C(1u,1v)1C(u,v)=u+v+C(1-u,1-v)-1 , for $(u,v)\in[0,1]^{2}$.
Proof.
The proof of (iii) is given by Nelsen, (2006) (Theorem 2.7.3) and (i) and (ii) follow by a similar method. ∎
This result implies that the random vector $(U,V)^{\top}=(F_{Y}(Y),F_{Z}(Z))^{\top}$, or its df $C$, inherits the properties of h-symmetry, v-symmetry, radial and joint symmetry from $(Y,Z)^{\top}$; the corresponding lines and centre of symmetry are at 0.5 and $(0.5,0.5)$. The next result provides the first connection between symmetries of copula functions and v-transforms. The function $\mathcal{V}(u)=|2u-1|$ appearing in this result is the simplest example of a v-transform.
Proposition 1.
Let $C$ be the distribution function of $(U,V)^{\top}$. Then, for all $(u,v)\in[0,1]^{2}$,
  1. (i)
    if $C$ is h-symmetric, $C(u,v)=(u+(-1)^{I_{\{v<0.5\}}}C^{(h)}(u,|2v-1|))/2$ where $C^{(h)}$ is the df of $(U,|2V-1|)^{\top}$;
  2. (ii)
    if $C$ is v-symmetric, $C(u,v)=(v+(-1)^{I_{\{u<0.5\}}}C^{(v)}(|2u-1|,v))/2$ where $C^{(v)}$ is the df of $(|2U-1|,V)^{\top}$;
  3. (iii)
    if $C$ is jointly symmetric, $C(u,v)=(2u+2v-1+(-1)^{I_{\{(u-0.5)(v-0.5)<0\}}}C^{*}(|2u-1|,|2v-1|))/4$ where $C^{*}$ is the df of $(|2U-1|,|2V-1|)^{\top}$.
Proof.
We prove (i) and (iii) since (ii) follows immediately from (i). We use the equalities $C^{(h)}(u,v)=\mathbb{P}(U\leqslant u,|2V-1|\leqslant v)=\mathbb{P}(U\leqslant u,|V-0.5|\leqslant v/2)=C(u,(1+v)/2)-C(u,(1-v)/2)$ and part (i) of Lemma 1 to infer that $C^{(h)}(u,|2v-1|)=C(u,1-v)-C(u,v)=u-2C(u,v)$ if $v<0.5$ and $C^{(h)}(u,|2v-1|)=C(u,v)-C(u,1-v)=2C(u,v)-u$ if $v\geqslant 0.5$. Hence part (i) follows.
To prove (iii) note that, if $C$ is jointly symmetric, then $C^{(h)}$ is v-symmetric since $C^{(h)}(u,v)+C^{(h)}(1-u,v)=C(u,(1+v)/2)-C(u,(1-v)/2)+C(1-u,(1+v)/2)-C(1-u,(1-v)/2)=(1+v)/2-(1-v)/2=v$. Hence $C^{(h)}(u,|2v-1|)=(|2v-1|+(-1)^{I_{\{u<0.5\}}}C^{*}(|2u-1|,|2v-1|))/2$. In combination with part (i) this yields part (iii).
If $C$ is the copula of the h-symmetric random vector $(Y,Z)^{\top}$ in part (i) of Lemma 1, $C^{(h)}$ in part (i) of Proposition 1 is the copula of $(Y,|Z-b|)^{\top}$. This follows by observing that if $V=F_{Z}(Z)$ is the probability-integral transform of $Z$, then the probability-integral transform of |Zb||Z-b| is $F_{|Z-b|}(|Z-b|)=|2V-1|$. Similarly, if $C$ is the copula of the jointly symmetric random vector $(Y,Z)^{\top}$, then $C^{*}$ in part (iii) of Proposition 1 is the copula of $(|Y-a|,|Z-b|)^{\top}$.
In this paper we apply the symmetry concepts to random vectors $(Y,Z)^{\top}$ with a joint density $f_{Y,Z}(y,z)>0$ on $\mathbb{R}^{2}$, implying that the copula $C$ has density $c$ given by
$c(u,v)=f_{Y,Z}(F_{Y}^{-1}(u),F_{Z}^{-1}(v))/\left(f_{Y}(F_{Y}^{-1}(u))f_{Z}(F_{Z}^{-1}(v))\right),$
where $f_{Y}$ and $f_{Z}$ denote the marginal densities. It is easily inferred from Definition 2 that h-symmetry of $(Y,Z)^{\top}$ is equivalent to $f_{Y,Z}(y,b+z)=f_{Y,Z}(y,b-z)$ for all $(y,z)\in\mathbb{R}^{2}$ and some $b\in\mathbb{R}$, while v-symmetry and radial symmetry are equivalent to $f_{Y,Z}(a+y,z)=f_{Y,Z}(a-y,z)$ for some $a\in\mathbb{R}$ and $f_{Y,Z}(a+y,b+z)=f_{Y,Z}(a-y,b-z)$ for some $a,b\in\mathbb{R}$. Similarly, h-symmetry of the copula $C$ is equivalent to the identity c(u,v)=c(u,1v)c(u,v)=c(u,1-v), v-symmetry is equivalent to c(u,v)=c(1u,v)c(u,v)=c(1-u,v) and radial symmetry is equivalent to c(u,v)=c(1u,1v)c(u,v)=c(1-u,1-v). When densities exist, the following Corollary of Proposition 1 is immediate.
Corollary 1.
Let $C$ be the distribution function of $(U,V)^{\top}$ and assume $C$ has joint density $c$. For all $(u,v)\in[0,1]^{2}$,
  1. (i)
    if $C$ is h-symmetric, $c(u,v)=c^{(h)}(u,|2v-1|)$ where $c^{(h)}$ is the density of $(U,|2V-1|)^{\top}$;
  2. (ii)
    if $C$ is jointly symmetric, $c(u,v)=c^{*}(|2u-1|,|2v-1|)$ where $c^{*}$ is the density of $(|2U-1|,|2V-1|)^{\top}$.
Simple calculations show that part (ii) also implies the following series of equations which will be used in the sequel:
$c^{*}(u,v)=c\left(\frac{1-u}{2},\frac{1-v}{2}\right)=c\left(\frac{1+u}{2},\frac{1-v}{2}\right)=c\left(\frac{1-u}{2},\frac{1+v}{2}\right)=c\left(\frac{1+u}{2},\frac{1+v}{2}\right).$ (4)
@@ 2.3 The first-order copula of GARCH(1,1)-type processes
We can now apply these symmetry ideas to GARCH-type processes. Simple calculations show that the joint density of a pair of successive variables $(X_{1},X_{2})^{\top}$ in a process of GARCH(1,1) type is given by
@@
where $f_{\sigma}$ is the marginal density of the volatility process $(\sigma_{t})_{t\in\mathbb{Z}}$ and $f_{\epsilon}$ is the density of the innovations. The density $c_{1}$ of the copula $C_{1}$ of $(X_{1},X_{2})^{\top}$ is thus
@@
The key insight from these equations is the following.
@@
    If the volatility function $\sigma(x,s)$ is an even function of $x$, this also implies that $f_{X_{1},X_{2}}(x,y)=f_{X_{1},X_{2}}(-x,y)$ and so $(X_{1},X_{2})^{\top}$ is v-symmetric about zero and hence radially and jointly symmetric about the origin.
In case (i) above we can apply Lemma 1 to infer that the copula $C_{1}$ is h-symmetric and we can apply Proposition 1 and Corollary 1 to infer that its density may be written in the form $c_{1}(u,v)=c_{1}^{(h)}(u,|2v-1|)$ where $c_{1}^{(h)}$ is the density of the copula of $(X_{1},|X_{2}|)^{\top}$. In case (ii) we infer that the copula $C_{1}$ is jointly symmetric and its density may be written in the form $c_{1}(u,v)=c_{1}^{*}(|2u-1|,|2v-1|)$ where $c_{1}^{*}$ is the density of the copula of $(|X_{1}|,|X_{2}|)^{\top}$. In the latter case it follows from (4) that $c_{1}^{*}(u,v)=c_{1}((u+1)/2,(v+1)/2)$, which means that the copula density $c_{1}^{*}$ may be viewed as a blown-up version of the upper quadrant of the copula density $c_{1}$, i.e. the part of the density on $0.5,1]^{2}$.
Despite the strong symmetry revealed by Proposition 2, the bivariate distribution of $(X_{1},X_{2})^{\top}$ in a symmetric process of GARCH(1,1) type is not in general exchangeable. This implies that ARCH and GARCH are not time-reversible models; in general $f_{X_{2}\mid X_{1}}(y\mid x)\neq f_{X_{1}\mid X_{2}}(y\mid x)$ and thus, in an intuitive sense, they are models in which the predictive distribution of the future given the past differs from the “predictive distribution” of the past given the future. From an econometric point of view, this would be considered natural. The non-exchangeability carries over to the copula so that we have in general that $C_{1}(u,v)\neq C_{1}(v,u)$ on $0,1]^{2}$.
@@ 2.4 Computational examples
In this section we use computational methods to approximate and plot the copula density $c_{1}$ in a number of models of GARCH(1,1) type; in Section S2 of the Online Supplement we give some accompanying pictures of the joint density $f_{X_{1},X_{2}}$ for selected models.
We concentrate on processes of ARCH(1) type in our illustrations since their first-order Markov structure simplifies computation of the unknown marginal density $f_{X}$ and distribution function $F_{X}$, which are key ingredients in (8). Beginning with the equation
@@
We also consider the classical GARCH(1,1) model with Gaussian innovations. To analyse this process we use simulation and estimation to find a parametric model for the unknown marginal density $f_{\sigma}$ of the volatility process. We can then integrate out volatility in expressions like (5), (6) and (7) [^2] All details of the model for $f_{\sigma}$ and the discretization method for ARCH(1)-type processes can be found in the code that accompanies this paper at https://github.com/ajmcneil/papers.. The parameter values for all the models we consider are summarized in Table 1. In the examples that follow we concentrate on illustrating the copulas densities $c_{1}$; corresponding illustrations of the joint densities $f_{X_{1},X_{2}}$ can be found in Section S2 of the Supplementary Material.
@@
Table 1: Parameter values for processes analysed in paper. The volatility function $\sigma(x,s)$ in all models satisfies $\sigma(x,s)^{2}=\alpha_{1}+(\alpha_{1}+\gamma_{1}I_{\{x<0\}})x^{2}+\beta_{1}s^{2}$. Innovation distributions used are either Gaussian, Student t with $\xi=4$ or $\xi=2.5$ degrees of freedom or skewed Student t with $\xi=4$ degrees of freedom and skewness parameter $\lambda=0.8$; all innovation distributions are scaled to have mean zero and variance one.
@@
2.4.1 Classical ARCH(1) process with Gaussian innovations (symmetric)
For the parameter values in Table 1, the contours of the copula density $c_{1}$ are shown in the top left panel of Figure 1 The non-exchangeability, evidenced by a lack of symmetry in the dashed line $v=u$, is evident. We can also observe the joint symmetry, that is the h-symmetry and v-symmetry also implying rotational symmetry of order 2. Changing $\alpha_{1}$ would change the degree of dependence; in particular, increasing $\alpha_{1}$ would strengthen the cross-shaped character of the copula density.
In Section S2 of the Supplementary Material we also include a contour plot of the absolute value copula for this model, that is the copula of $(|X_{1}|,|X_{2}|)^{\top}$. The picture for the absolute value copula is confirmed to be a blown-up version of the top right quadrant of the contour plot of the copula density $c_{1}$.
2.4.2 Classical ARCH(1) process with Student innovations (symmetric)
@@
This is shown in the top right panel of Figure 1 for a process with the same parameters as in the previous example and degree of freedom $\xi=4$ for the innovation distribution. Remarkably, the copula appears to be very close to an exchangeable copula for this combination of parameters; the joint symmetry is again apparent. The bottom left panel of Figure 1 (see also Figure S1) show a process in which $\xi=2.5$ which shows non-exchangeability more clearly.
@@
It is notable that, although the heaviness of the tail of the marginal distribution increases with decreasing degree of freedom, the dependence actually weakens. We can estimate a distance $D_{1}=\int_{0}^{1}\int_{0}^{1}|c_{1}(u,v)-1|\mathrm{d}u\mathrm{d}v$ between the copula density $c_{1}$ and the density of the independence copula; for the Gaussian, t4 and t2.5 copulas we obtain respectively 0.213, 0.158 and 0.092.
2.4.3 Classical GARCH(1,1) process with Gaussian innovations (symmetric)